OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (amistre64):

1 = 4m+b 2 = 5m+b -1 = -4m-b 2 = 5m+b ------------ 1 = m ; b= -3

OpenStudy (amistre64):

-x + y = -3\ and since they dont like negative first terms ..... x - y = 3

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

would you be able to do it with other points?

OpenStudy (anonymous):

i think so

OpenStudy (amistre64):

my method was different than what the books usually teach at this stage. I used a system of equations to determine the values of the slope and y intercept

OpenStudy (anonymous):

oh okay i see what you mean

OpenStudy (amistre64):

good luck